OpenStudy (anonymous):

@AnswerMyQuestions @$w3G_Godd

OpenStudy (anonymous):

A circle has a diameter of ten millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth. fifteen point seven millimeters; seventy-eight point five square millimeters thirty-one point four millimeters; seventy-eight point five square millimeters fifteen point seven millimeters; three hundred fourteen point zero square millimeters thirty-one point four millimeters; three hundred fourteen point zero square millimeters

OpenStudy (anonymous):

given: diameter = 10mm 2* radius = diameter radius = diameter/2 radius = 10/2 radius = 5 radius r is 5mm

OpenStudy (anonymous):

so whitch one @$w3G_Godd

OpenStudy (anonymous):

Circumference C = 2(pi)r C = 2(3.14)5 C = 10(3.14) C = 31.4mm Area A = (pi)r^2 A = (3.14)(5)^2 A = (3.14)(25) A = 78.5 square mm Sorry i never finshed

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

A square pyramid has a total surface area of one thousand square feet. The area of the base is two hundred square feet. What is the area of one triangular face of the lateral surface? one hundred twenty-five square feet one hundred fifty square feet one hundred seventy-five square feet two hundred square feet

OpenStudy (anonymous):

bs=400 Area of a triangle is (1/2)bs=(1/2)*400=200